# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v1

@@ 5 Native loss detection from logical teleportation
Figure 7: Deep logical algorithms with teleported gates. (a) We study a deep circuit consisting of teleported $X$ and $Z$ logical gates, similar in structure to small-angle synthesis algorithms [34, 35]. (b) Circuit-level simulation results for 11 layers of random teleported $Z$ and $X$ logical gates, with all logical qubits initialized in the presence of noise (physical error rate $p=1\%$ and $d=7$). The delayed-erasure decoder (pink) substantially outperforms an MLE decoder that does not account for loss (black) and improves with increasing loss fraction. The erasure channel (yellow) provides a lower bound.
As a proof of concept, we consider the teleportation-based logical circuit shown in Fig. 7(a), which mirrors the structure of the small-angle synthesis algorithm [34, 35]. This algorithm constructs small-angle rotations through sequences of $H$ and $T$ gates, with the latter often realized via teleportation. Consequently, loss detection is integrated into the process of executing the logical algorithm. The logical error rates for the circuit in Fig. 7(a) are plotted in Fig. 7(b) as a function of the loss fraction using two decoders: a delayed-erasure decoder leveraging SSR, and a regular MLE decoder. The delayed-erasure decoder demonstrates significant performance improvements, with algorithmic logical error rates rapidly decreasing as the loss fraction increases, achieving performance close to the lower bound of an erasure channel. Notably, this circuit is similar to the teleportation-based SE method investigated here in the memory setting, both utilizing SSR and replacing atoms frequently through teleportation without additional overhead. These results highlight the advantages of exploiting native loss detection within teleportation gadgets in logical circuits while also utilizing loss information to improve decoding. Finally, while our simulations focus on Clifford gate teleportation, they are expected to extend to non-Clifford gates, as the underlying loss detection principles remain consistent.
@@ Acknowledgments
SE method Average lifecycle length of data qubits Average lifecycle length of measure qubits Threshold for $L=1$ [%] Space overhead Time overhead Extra requirements
Conventional $4d$ 4 None $\sim 2d^{2}-1$ $\sim 4d$ None
SWAP period 1 8 8 $2.0\pm 0.08$ $\sim 2d^{2}-1$ $\sim 4d$ SSR
Free period 2 8 4 $3.63\pm 0.06$ $\sim 2d^{2}-1$ $\sim 4d$ Erasure conversion [17, 10]
Teleportation-based 4 4 $4.61\pm 0.05$ $\sim 3d^{2}-1$ $\sim 6d$ SSR
Free period 1 4 4 $4.68\pm 0.08$ $\sim 2d^{2}-1$ $\sim 4d-1$ Erasure conversion [17, 10]
Free period 0.25 (erasure channel) 1 1 $6.93\pm 0.1$ $\sim 2d^{2}-1$ $\sim 4d-1$ Erasure conversion [17, 10]
Table 1: Average lifecycles, error thresholds, and space-time overheads for different SE methods over $d$ repetitive SE rounds using the surface code $[[d^{2},1,d]]$. The relationship between lifecycles and logical error thresholds is highlighted in loss-dominated regimes ($L=1$).

### v2

@@ 5 Native loss detection from logical teleportation
Figure 7: Deep logical algorithms with teleported gates. (a) We study a deep circuit consisting of teleported $X$ and $Z$ logical gates, similar in structure to small-angle synthesis algorithms [35, 36]. (b) Circuit-level simulation results for 11 layers of random teleported $Z$ and $X$ logical gates, with all logical qubits initialized in the presence of noise (physical error rate $p=1\%$ and $d=7$). The delayed-erasure decoder (pink) substantially outperforms an MLE decoder that does not account for loss (black) and improves with increasing loss fraction. It also matches the performance of a decoder with perfect loss information (gray, dashed), and for loss fractions less than 1, approaches that of the erasure channel (khaki, dashed). The erasure channel performance is achieved using Free SE with period 0.25, corresponding to loss detection and qubit replacement after every gate, thus providing a lower bound.
As a proof of concept, we consider the teleportation-based logical circuit shown in Fig. 7(a), which mirrors the structure of the small-angle synthesis algorithm [35, 36]. This algorithm constructs small-angle rotations through sequences of $H$ and $T$ gates, with the latter often realized via teleportation. Consequently, loss detection and correction are integrated into the process of executing the logical algorithm. The logical error rates for the circuit in Fig. 7(a) are plotted in Fig. 7(b) as a function of the loss fraction using various decoders: a delayed-erasure decoder leveraging SSR, a regular MLE decoder, and an MLE decoder with perfect loss location information. The delayed-erasure decoder demonstrates significant performance improvements, with algorithmic logical error rates rapidly decreasing as the loss fraction increases, achieving the performance of the decoder with perfect loss information. Furthermore, for loss fractions less than 1, it achieves the lower bound set by the erasure channel—corresponding to loss detection and qubit replacement after every gate. Notably, this circuit is similar to the teleportation-based SE method investigated here in the memory setting, both utilizing SSR and replacing atoms frequently through teleportation without additional overhead. These results highlight the advantages of exploiting native loss detection within teleportation gadgets in logical circuits while also utilizing loss information to improve decoding. Finally, while our simulations focus on Clifford gate teleportation, they are expected to extend to non-Clifford gates, as the underlying loss detection principles remain consistent.
@@ 6 Conclusion
Note added. Since the completion of our work, we became aware of several related studies examining the effect of qubit loss on a single logical memory: teleportation-SE [22], loss detection units [23], and SWAP-SE [76]. While addressing similar challenges, these works differ from ours in focus and methodology. Notably, improved thresholds have been recently reported in updated versions of Refs.[22, 76], based on the use of decoding strategies similar to the delayed erasure decoder described in this work. These contributions also utilize error models slightly different from those presented here. For completeness, our updated Appendix A includes results obtained using the error models utilized in Refs.  [22, 76].
@@ Appendix A SE Methods Lifecycles and Space-time Overheads
Category Conventional SWAP period 1 Free period 2 Teleportation-based Free period 1 Free period 1 + exact loss info Free period 0.25 + exact loss info
Average lifecycle length of data qubits $4d$ 8 8 4 4 1 (detection) / 4 (replacement) 1 (detection) / 1 (replacement)
Average lifecycle length of measure qubits 4 8 4 4 4 1 (detection) / 4 (replacement) 1 (detection) / 1 (replacement)
Space overhead $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 3d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$
Time overhead $\sim 4d$ $\sim 4d$ $\sim 4d$ $\sim 6d$ $\sim 4d-1$ $\sim 4d-1$ $\sim 4d-1$
Extra experimental requirements None SSR Erasure conversion [17, 10], detection and replacement every 2 SE rounds SSR Erasure conversion [17, 10], detection and replacement every SE round Erasure conversion [17, 10], detection every gate, and replacement every SE round Erasure conversion [17, 10], detection and replacement every gate
Threshold for $L=1$ [%] (Channel 1) NA $2.27\pm 0.02$ $3.98\pm 0.06$ $4.78\pm 0.01$ $5.07\pm 0.02$ $7.16\pm 0.01$ $9.51\pm 0.04$
Effective distance for $d=7$ (Channel 1) NA $6.41\pm 0.27$ $6.77\pm 0.2$ $7.19\pm 0.45$ $6.83\pm 0.37$ $6.97\pm 0.24$ $7.46\pm 0.36$
Threshold for $L=1$ [%] (Channel 2) NA $1.94\pm 0.01$ $2.75\pm 0.07$ $3.5\pm 0.05$ $3.76\pm 0.01$ $5.12\pm 0.01$ $6.5\pm 0.01$
Effective distance for $d=7$ (Channel 2) NA $6.13\pm 0.26$ $5.5\pm 0.23$ $6.5\pm 0.21$ $6.32\pm 0.87$ $6.45\pm 0.47$ $7.47\pm 0.49$
Table 1: Comparison of SE methods under various error models. Each column represents a method, and each row describes a property such as average qubit lifecycles, thresholds, effective distance and overheads. The final column estimates performance under very frequent detection and replacement, providing the erasure channel bound. Channel definitions for $L=1$ are as follows. Channel 1: $p^{\prime}\{L\otimes I,I\otimes L\}$. Channel 2: $\frac{p^{\prime}}{2}\{L\otimes Z,Z\otimes L,L\otimes I,I\otimes L\}$. $p^{\prime}=1-\sqrt{1-p}\sim p/2$ where $p$ is the entangling gate physical error rate.
@@
Appendix B Delayed-Erasure Decoder
In this section, we will describe our approach for decoding in the presence of loss. We first describe the exact MLE solution. Then, we describe the approximated MLE decoding which we use in the main text.
B.1 Exact MLE Decoding in the Presence of Loss
In this section, we provide a formulation of the most-likely-error (MLE) decoding problem in the presence of loss errors. Loss acts in a much more complex fashion compared to normal Pauli errors, and therefore require more care in terms of their formulation. The results also depend on the details of the error model and measurement model, introducing further complications. We will attempt to describe the full formulation in this section, which will then guide the choice of heuristic methods, primarily based on reweighting edges in the decoding graph, that will approximate the MLE problem.
We formulate the problem by generalizing the discussion in Ref. [12]. We can associate a binary variable $E_{i}$ for each Pauli error, and similarly $L_{i}$ for each loss event. We assume that the loss and Pauli errors are applied independently (i.e. the channel does not simultaneously apply two types of errors, even though that could be the case when doing a CZ gate, where there is a correlation between loss on one qubit and a Pauli error on the other). When running the circuit, we obtain information about errors and loss through qubit measurements. This information comes in two flavors: First, there are flags $F_{j}$, which indicate a qubit was lost when we tried to measure it. Second, there are detectors $D_{j}$, which are products of stabilizer measurement results that should be deterministic in the absence of errors. A non-trivial aspect of loss decoding is that the detectors must be determined based on the flags, since some qubit measurement results are labeled as “invalid” instead of 0 or 1.
The most likely error problem seeks to find the assignment of Pauli errors $\vec{E}$ and loss errors $\vec{L}$ that maximizes the following quantity:
$P(\vec{E},\vec{L}|\vec{F},\vec{D}),$ (1)
Namely, given the observed flags, and the resulting detectors that we define conditional on seeing those flags, what is the most likely error.
We can rewrite this expression using Bayes’ rule
$P(\vec{E},\vec{L}|\vec{F},\vec{D})$ $=\frac{P(\vec{E},\vec{L},\vec{F},\vec{D})}{P(\vec{F},\vec{D})}=\frac{P(\vec{F},\vec{D}|\vec{E},\vec{L})P(\vec{E},\vec{L})}{P(\vec{F},\vec{D})}.$ (2)
Since we are given a fixed observation $\vec{F}$, $\vec{D}$, the denominator is fixed. We therefore seek to maximize the product of the probability of the error configuration $P(\vec{E},\vec{L})$ and the conditional probability of observing these syndromes. In the presence of loss, an important distinction from the usual case is that the latter factor is no longer always 0 or 1.
The error probability $P(\vec{E},\vec{L})$ can be readily written down based on the log likelihood ratios of individual events. One potential subtlety that one may need to worry about is that for a given qubit at a given lifecycle, only one loss event can occur. Thus, if we have a sequence of gates with loss probability $p_{1}$, $p_{2}$, …, then the loss event probabilities may need to be written as $p_{1}$, $(1-p_{1})p_{2}$, $(1-p_{1})(1-p_{2})p_{3}$,…, although the deviations may be higher-order in practice.
The more challenging factor to evaluate is $P(\vec{F},\vec{D}|\vec{E},\vec{L})$. First, we need to specify how the new detectors are chosen given a heralded loss pattern. In Ref. [42], they compare the stabilizer value against the next available syndrome measurement result at the same site. This seems like a reasonable strategy, since this bit of information is really not accessible until the next time we measure this stabilizer, and is not revealed by any other stabilizers. Note that strictly speaking, if we solve the full MLE problem, the precise choice of detector shouldn’t matter due to stabilizer equivalences, so long as it forms a complete basis. However, it will matter if we are using heuristic decoders. After forming the detectors, we need to evaluate whether a given loss and error pattern is consistent with the observed loss flags and detectors. The loss flag check is straightforward, and can be imposed as a constraint that exactly one of the loss locations corresponding to a given qubit are activated. The detector is more subtle, because now some of the way they are triggered becomes probabilistic. A brute force way to check would be as follows: For a given loss pattern, the circuit is now known. We can forward propagate all error events through the new circuit, and obtain measurement results, which will then tell us the probability distribution and allow us to calculate the conditional probability. Unfortunately, this seems rather costly to perform and may not be as easily phrased as a simple optimization problem with a linear optimization target, so we need to think a bit more about better ways to evaluate this probability.
Finally, we comment on the benefits and downsides of heuristic edge re-weighting techniques. For example, in Ref. [42], given a leakage event, they assign an $X$ and $Z$ error on the $i$th location with probability $i/2n$ and propagate this error forward to modify the edge weights. Note that the detailed error event is related to their leakage model, where all qubits that interact with a leaked qubit experience large depolarizing noise. This is good to leading order, since it captures the fact that qubits that interact with the measure qubit later are more likely to have experienced leakage. However, it seems to still miss one potentially important aspect: If a qubit is leaked at some point, then it is leaked for the rest, so the error events are actually correlated. This correlation can in principle be leveraged to more accurately deduce the error that occurred.
Given these considerations, the main goal is thus to find good heuristic methods that approximate MLE. In particular, some better intuition about the types of error events that are incorrectly classified for a given weighting could help inform modifications that make it closer to MLE. One could also either try different heuristics and see what gives the best logical error rate (focusing perhaps on “hard” instances to make the search more efficient), or implement MLE so that it also becomes clear how big the optimality gap to MLE is for a given heuristic (this also allows one to more clearly focus on concrete instances of heuristic decoding failures). With the possible exception of neural network decoders that automatically handle both aspects (see Google’s paper, where there are actually also simulations with heralded leakage [5]), it seems overall that closing the gap to MLE may make a bigger difference than the MLE-MLD gap.
B.2 Approximate MLE Decoding
In this section, we describe our approach for constructing a decoder to handle losses by approximating the MLE decoding introduced in the previous section.
We utilize the concept of qubit lifecycles: the number of potential locations where a physical qubit may be lost—due to single-qubit gates, multi-qubit gates, initialization, idling, measurement, or other processes. Given a heralded loss upon measurement, the decoder identifies a loss occurring in one of these locations, each characterized by its specific probability of loss.
Consider a scenario where, at the end of qubit $i$’s lifecycle $C_{i}$, a loss is detected, with $L_{ij}$ representing potential loss events, each associated with a probability $p_{ij}$. For each potential loss event $L_{ij}$, we simulate the effect of the loss by constructing a loss circuit. This involves removing subsequent gates acting on the lost qubit and modeling the measurement of the lost qubit as random. Using Stim’s gauge detector infrastructure [53], we generate the decoding graph for each specific loss event, denoted as $\text{DEM}_{ij}$.
The decoding graph is a collection of hyperedges and their associated probabilities. Each error mechanism corresponds to a hyperedge that connects the detectors it triggers. Detectors are defined as the products of stabilizer measurements and are expected to be $+1$ in the absence of errors.
Each decoding graph $\text{DEM}_{ij}$ is a collection of error probabilities $p_{n}$ and corresponding detector configurations $D_{n}$:
$\text{DEM}_{ij}=\{(D_{n},p_{n})\},$ (3)
where $D_{n}=[D_{n}(0),D_{n}(1),\cdots]$ represents the list of detectors triggered by the error event. Repeating this process for all potential loss events $L_{ij}$ within a lifecycle $C_{i}$ results in a set of decoding graphs $\text{DEM}_{ij}$ and their associated probabilities $p_{ij}$.
To compute the final decoding graph for a lifecycle $C_{i}$, we sum over all decoding graphs for the potential loss events:
$\text{DEM}_{i}=\sum_{j}p_{ij}\cdot\text{DEM}_{ij}.$ (4)
This summation is straightforward because only one of these events can occur.
When multiple qubit losses occur in different lifecycles, loss errors can cause correlated Clifford errors. Combining syndromes from distinct lifecycles may interfere non-linearly, resulting in syndromes that cannot be derived from independent consideration of loss events. While the optimal decoder would consider all combinations of potential loss locations for all qubits, this approach scales poorly with multiple lossy lifecycles and potential loss events, making it impractical for small code distances ($d<7$).
To address this, we explore a decoder that evaluates each lifecycle independently and averages over them. Additionally, we introduce the option to include a single combination of losses across different lifecycles. Specifically, we consider the first potential loss locations $L_{i1}$ for all lossy lifecycles $L_{i}$ and generate a corresponding decoding graph, $\text{DEM}_{\text{first comb}}$.
The decoding graphs for individual potential loss locations $L_{ij}$ are generated in a pre-processing step, independent of any specific error model or probabilities of Pauli and loss errors. During real-time decoding, these graphs are summed using the specific error model probabilities $p_{ij}$. Note that we separately account for loss errors and Pauli errors. Lossless circuits (without losses) are used to efficiently generate the decoding graph for Pauli errors, denoted $\text{DEM}_{\text{Pauli}}$.
For a specific shot with a heralded loss pattern using SSR and a given error model, the final decoding graph is computed by summing: 1. Lossy lifecycle decoding graphs $\text{DEM}_{i}$, 2. The Pauli decoding graph $\text{DEM}_{\text{Pauli}}$, and 3. The first potential loss combination decoding graph $\text{DEM}_{\text{first comb}}$.
The final decoding graph is given by:
$\text{DEM}_{\text{final}}=\sum_{i}\text{DEM}_{i}+\text{DEM}_{\text{Pauli}}+\omega\cdot\text{DEM}_{\text{first comb}},$ (5)
where $\omega$ is a combination weight determined based on the analysis in Appendix Section B.2.B.2.1.
The summation of probabilities is calculated using the following equation:
$\sum_{i=1}^{l}p_{i}\prod_{j\neq i}(1-p_{j})+\sum_{i=1}^{l}\sum_{j=i+1}^{l}\sum_{k=j+1}^{l}p_{i}p_{j}p_{k}\prod_{m\neq i,j,k}(1-p_{m}).$ (6)
This ensures accuracy up to $O(p^{3})$.
Finally, given the final decoding graph, any known decoder can be used to decode the data for a specific shot, such as MWPM, MLE, or other methods.
B.2.1 Adjusting Loss Combination Weight
In this section, we explore the influence of the weight of the combination, $\omega$, on the efficacy of the delayed erasure decoder, which considers both independent error events and the first combination of potential loss events. As described earlier, the decoder’s decision-making process involves summing different decoding graphs (see Eq. 5). To incorporate the impact of individual loss events, the first loss combination event is scaled by $\omega$ along with the standard Pauli error decoding graph.
To benchmark the performance of our decoder and determine the optimal value of $\omega$, we performed circuit-level simulations on a memory logical qubit with multiple rounds of conventional syndrome extraction (SE). We varied $\omega$ between 0 and 1. For $\omega=0$, each lifecycle’s decoding graph is considered independently, without including an additional decoding graph for the first combination.
The results for the conventional SE method are presented in Fig. 8. Non-zero values of $\omega$ degrade the decoder’s performance. However, it is important to note that these results pertain to the conventional SE method, and different SE methods employing distinct gate sets may behave differently. For simplicity, the numerical simulations presented in this paper assume $\omega=0$.
Figure 8: Logical memory simulations with $d$ rounds of the conventional SE method. Variation of the logical error rate as a function of the combination weight $\omega$ for distance 9, across different physical error rates ($p=0.01$ and $p=0.03$). The results indicate that nonzero values of $\omega$ deteriorate the decoder’s performance.